[Search for users] [Overall Top Noters] [List of all Conferences] [Download this site]

Conference asdg::cmp

Title:CMP
Notice:Welcome to the CMP notesfile
Moderator:FABSIX::B_TOWERLLE
Created:Tue Oct 31 1995
Last Modified:Fri Jun 06 1997
Last Successful Update:Fri Jun 06 1997
Number of topics:94
Total number of notes:6381

47.0. "CMP-SCRUB.A2 EE/PE PASSDOWNS" by SUBPAC::BJUBINVILLE (NFD Fire - Rescue) Sat Dec 09 1995 00:26

    
        THIS NOTE IS DESIGNATED FOR THE EQUIPMENT AND PROCESS
      ENGINEERING PASSDOWN FOR:
    
                                   CMP-SCRUB.A2
T.RTitleUserPersonal
Name
DateLines
47.59Inside brush station leakingPATE::GEEMon Nov 06 1995 06:275
	* The inside brush station was leaking DI water.

	Replaced the front roller shaft o-ring, passed mqc's.

					B shift
47.60robot fingers adjustedFABSIX::R_GEETue Nov 21 1995 07:159
	* The robot wouldn't pickup wafers out of the spin station.

	  - Checked for any loose parts, none found.

	  - Had to adjust the robot finger offset from 310 to 480.

	We will need to keep an eye on this, the position may be drifting.

					B Shist
47.1INSTALLATION AND MECHANICAL ACCEPTANCESUBPAC::BJUBINVILLENFD Fire - RescueFri Dec 08 1995 23:49173
SUBPAC::LANDRY                                       13 lines   6-JUN-1994 17:02
                                -< 6/6 update >-
--------------------------------------------------------------------------------
    All,
    
    KSI has been working on hooking up the exhaust/DI water drain.  All
    seemed fine at noon time, with the plumbing attached to the system
    (but unglued).  After my new/improved protocol class, I went back into
    the fab to see where everything stood.  Much to my surprise, the
    plumbing from the back of the system had been removed and was lying
    on the floor next to the system.  There wasn't anyone around to answer
    why this is (maybe it's off to be PVC cemented???).  I'll follow up on
    this in the morning.  No other activity seems to have taken place with
    this system.
    
                                                  Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 66.2                         CMP-SCRUB.A2                          2 of 125
SUBPAC::LANDRY                                       19 lines  16-JUN-1994 16:10
                                -< 6/16 Update >-
--------------------------------------------------------------------------------
    Installation Update:
    
    The electrical is complete, including the OCP.  The OCP still needs the
    communication line tie in.
    
    The Ammonium Hydroxide line has been installed.
    
    The high purity DI line has been run to the rack above the floor. 
    Still needs to be tied into the main system.
    
    The GCW drain section has been installed.  Still needs to be tied into
    the exhaust duct, which has yet been run.
    
    Earl Dawson and I moved all the accessory equipment up to the fab and
    installed the "touch screen" monitor and one of the smoked grey covers.
    There is still one more cover to attach at this time.  Need hardware to
    accomplish the feat.
    
                                                Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 66.3                         CMP-SCRUB.A2                          3 of 125
SUBPAC::TOOKER                                       24 lines   8-JUL-1994 11:04
                         -< Hook-Up Status For 7/8/94 >-
--------------------------------------------------------------------------------
                CMP_SCRUB.A2 Tool Hook-Up Status For 7/8/94



COMPLETED ITEMS:
----------------

    *  DI Water connections

    *  Ammonium Hydroxide connections

    *  All GCW drain connections

    *  Eelectrical connections


OPEN ITEMS:
-----------

    *  CDA final connection

    *  All exhaust work

    *  OCP interface to system

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 66.4                         CMP-SCRUB.A2                          4 of 125
SUBPAC::TOOKER                                       14 lines  26-JUL-1994 09:22
                               -< CMP-SCRUB.A2 >-
--------------------------------------------------------------------------------

                CMP-SCRUB.A2 Tool Hook-Up Status As Of  7/25/94


PUNCH LIST ITEMS
----------------

             ITEM                                      RESPONSIBLE PERSON
   --------------------------------------------------------------------------

  1) Complete exhaust work.                            D. Atkinson/NEEVCO

  2) Install additional pipe                           D. Atkinson/NEEVCO
     support on exhaust.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 66.5                         CMP-SCRUB.A2                          5 of 125
SUBPAC::LANDRY                                        5 lines   1-AUG-1994 13:17
                     -< 8/1 equipment installation update >-
--------------------------------------------------------------------------------
    Brian and I hooked up the last plumbing connection (CDA).  The only
    remaining hookup is the communications connection between the main
    system and the OCP.
    
    							Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 66.6                         CMP-SCRUB.A2                          6 of 125
SUBPAC::BJUBINVILLE                                  16 lines   2-AUG-1994 16:44
                         -< Ron Simmons has arrived!! >-
--------------------------------------------------------------------------------
    
    
    
        Ron Simmons was in this afternoon so here's the scoop.
    The facility guys were turning on the N2, DI water and CDA when
    they discovered some leaks near CMP.A1 so the scrubbers were
    on hold.  Ron will be back in the morning around 8am to start
    things into purge.  The power is all hooked up but the 15A
    breaker is off on both machines.  There was a concern with a
    fitting on the 1/2" ammonium line.  Ron Simmons would rather
    not use stainless.  He doesn't need the fitting to put the
    machine into purge so he'll bring the fitting/parts with him
    on Monday.
    
    
    BJ

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 66.7                         CMP-SCRUB.A2                          7 of 125
SUBPAC::BJUBINVILLE                                  11 lines   4-AUG-1994 17:46
                            -< passdown for 8/4/94 >-
--------------------------------------------------------------------------------
    
        Doc and I got around 30psi for DI pressure for now.
    Bob Kerwin is looking into trying to get it up to 45psi.
    There was a leak detected later on in the afternoon by
    Larry.  Is was a slow drip but it eventually dispersed 
    a good size puddle.  Greg found the leak ands we seem
    to have stopped it.  We put the machine back into purge 
    until Monday.  Right now there seems to be a sensor problem
    within the spin chamber.  I'll be calling Ron Simmons in
    the morning regarding this.
                                                   BJ

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 66.10                        CMP-SCRUB.A2                         10 of 125
SUBPAC::LANDRY                                       12 lines  11-AUG-1994 19:10
                          -< purged and purified!!! >-
--------------------------------------------------------------------------------
    The H20 Purge on the ammonium Hydroxide line was completed at 3:00 PM.  
    The N2 purge on the line was then started and completed.  The NH4OH
    line was then re-attached to it's proper port.  We could possibly see
    NH4OH to the tool as soon as Friday.
    
    We still need to have the OCP's interfaced to the main system
    
    We attempted to run the systems today with a "burn-in" recipe.  The
    system expirienced several timeout/sensor problems.  The rep will
    be in to address the systems on Friday.
    
                                                    Doc
47.2MECHANICAL ACCEPTANCESUBPAC::BJUBINVILLENFD Fire - RescueFri Dec 08 1995 23:5011
SUBPAC::BJUBINVILLE                                   8 lines   9-AUG-1994 16:04
                       -< we're getting there here too! >-
--------------------------------------------------------------------------------
    
    
        Ron I and worked with the wafers on this scrubber
    today.  We installed the brushes and we cycled about (50)
    wafers without any problems.  We are going to purge 
    the NH4OH with DI in the morning.   All of the leaks
    have be found and repaired.
                                                   BJ
47.3SYSTEM NOT INITIALIZING OR HOMINGSUBPAC::BJUBINVILLENFD Fire - RescueFri Dec 08 1995 23:5140
SUBPAC::BJUBINVILLE                                  37 lines  15-AUG-1994 18:09
                             -< Passdown for 8/15 >-
--------------------------------------------------------------------------------
      
    
       A few problems this morning:    
    
    
    *  Ron and I came into where the SCRUB_A2 was stopped and had
       homing problems.  After we rebooted the machine, there was 
       a problem on boot up.  The machine is suppose to go right 
       into the OnTrak self initializing mode but it was staying 
       at the WinShell prompt.  After Ron called the software
       guru from OnTrack he had use copy a file over from another
       directory.  These machines are touch screens so we had to hunt
       down a keyboard with a RS232 connector.  We had to type the 
       following as these machines are not DOS operated.
    
         WinShell:> COPY DATA2 DATA -R
    
       After this is typed, reboot the machine it ok.  Every 2 minutes
       the machine does a self back-up of essential data files that usually
       takes a couple of milliseconds.  If you get acurate enough, if you
       shut the machine the exact same time as this back-up starts, the
       machine will put all of the robot offsets reset to factory defaults.
       OnTrack said this has only happened twice in our model history.  
       This is very rare.  I copied the offsets for both machines into
       their respective logbooks.  I will eventually copy them into here.
    
    *  The DI water is still an issue.  I spoke with Bob Kerwin and
       the issude is still being addressed.  He's looking at least
       another week before we some action.
    
    *  Repaired a few more leaks
    
    * we went over some maintainence training as well as some process
      issues with Steve.
    
    BJ
        
47.4brush motor speed out of specSUBPAC::BJUBINVILLENFD Fire - RescueFri Dec 08 1995 23:5113
SUBPAC::BJUBINVILLE                                  10 lines  17-AUG-1994 17:21
                               -< passdown 8/17 >-
--------------------------------------------------------------------------------
    
        A "inside brush motor speed out of spec" pops up now and then.  
    We traced it to the drive belts on the rollers.  We tightened them
    up and the motor feed back was stable.  The door switch in the 
    receiver tower was a giving both LOCKED and UNLOCKED at the same
    time if you pushed on the door.  The switch just needed some adjusting.
    Everything else is running great.
    
                                                        BJ
    
47.5BRUSH HEIGHTS, MOTOR SPEEDS ADJUSTEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Dec 08 1995 23:5217
SUBPAC::BJUBINVILLE                                  14 lines  18-AUG-1994 16:24
                             -< Passdown for 8/18 >-
--------------------------------------------------------------------------------
    
    
    *   Ron and I set the brush height to 14.5mm.  After some
     Oxide wafers were run on the Tencor, we found 14mm was the
     best height.
    
    *  We tweaked in the outside, inside and spin motor speed
     offsets for optimum stability.
    
    *  We've ran close to 600 wafers without any problems.  We
     should clear 1000 wafer for the mechanical acceptance
     tomorrow.
    
    BJ
47.6OCP INTERFACE CONFIG, PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Dec 08 1995 23:5222
SUBPAC::BJUBINVILLE                                  19 lines  22-AUG-1994 17:03
                             -< Passdown for 8/22 >-
--------------------------------------------------------------------------------
         This machine has been cycling wafers all day without any
    problems. 
         
         The OCP interface that was designed for us through OnTrak
    seems to be wrong.  Their schematic seems to have been misplaced
    so another one is being faxed to us tomorrow.  We will be evaluating
    this further tomorrow.  Jim T. will be looking the original specs
    that DEC layed out to OnTrak for our request.
    
         The ammonia needed to be regulated at 35psi. but it was never
    done.  Dan Smith is looking to get me a teflon regulator and teflon
    gauge in the morning.  Unregulated, it will be around 45-50psi.
    
         The DI water pressure was still sitting around 30psi at 3:30pm
    today.  Bob Kerwin did come into CMP to look at the gauge but it 
    remains to be unseen if anything was accomplished beyond 3:30pm.
    
    
    BJ
47.7ROLLER SHAFT LEAKSSUBPAC::BJUBINVILLENFD Fire - RescueFri Dec 08 1995 23:5321
SUBPAC::BJUBINVILLE                                  18 lines  25-AUG-1994 17:13
                             -< Passdown for 8/24 >-
--------------------------------------------------------------------------------
    
    
      *  Inner brush front roller shaft was leaking, replaced
         seal.  Note:  upon removal of the seal, it was discovered
         that the manufacturer installed the wrong seal.  Ron is
         aware of the seal and will speak to the main office about
         it.
    
      *  The inner exit manifold leaking out end (brush lift end),
         replaced manifold.

      *  The outer exit manifold leaking out end (brush lift end),
         replaced manifold.
    
      *  tweaked in inside, outside and spin motor speed offsets
    
    
    BJ
47.8TOUCH SCREEN FROZESUBPAC::BJUBINVILLENFD Fire - RescueFri Dec 08 1995 23:5319
SUBPAC::BJUBINVILLE                                  15 lines  29-AUG-1994 16:40
                             -< Passdown for 8/29 >-
--------------------------------------------------------------------------------
    
          The brush height was adjusted again for
    optimum cleaning.  
    
          The machine froze twice as a result to using
    the touch screem.  The new software might be here
    tomorrow.  This should take of problems like this.
    
          We still didn't receive any info on the machine
    tie in to the OCP.  With no ammonia, light process
    testing was done.  OnTrak has to confirmed that on
    two occasions, they faxed the tie in schematic to
    B+V, our previous riggers.  So, who knows what happened
    to it.  Once we get the ammonia on, process acceptance
    will begin. 
          
47.9SOFTWARE BACKED UPSUBPAC::BJUBINVILLENFD Fire - RescueFri Dec 08 1995 23:5411
SUBPAC::BJUBINVILLE                                   8 lines  31-AUG-1994 16:13
                             -< Passdown for 8/31 >-
--------------------------------------------------------------------------------
    
    
      Backed up entire drive and installed new software.  
    So far so good.  All of the bugs that we had are now gone.
    
      Still no ammonia, still waiting....
    
    BJ
47.10AMMONIA IS FLOWINGSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0116
SUBPAC::BJUBINVILLE "The other side of life -- 3rd"  13 lines  15-SEP-1994 16:27
                             -< Passdown for 9/15 >-
--------------------------------------------------------------------------------
    
    
         We got our ammonia finally turned on at 11am.  I've
    completed running the 150 wafers with the ammonia as
    requested by Steve.  There is still a IC chip that needs
    to be replaced on both OCP's.  This chip was suppose to
    be in either today or tomorrow.  Dan Smith will be taking
    care of this.  If you want the ammonia shut off for any
    reason, either hit the EMO button of the machine or 
    preferably the OCP.  I will fill Jim in on what Dan
    is going to do.
    
    BJ
47.11PROCESS ACCEPTANCE AND BRUSH HEIGHTSSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0120
ASDG::COOPERMAN                                      17 lines  23-SEP-1994 13:40
                          -< scrub.a2 passdown 9/22 >-
--------------------------------------------------------------------------------
    
    Jeff Farber came in on 9/22 to run some contamination wafers.
    During the visit the brush heights were readjusted different 
    from original setting back on 8/29 (ib was at about same height
    as ob). New settings are:
    
        	outside brush		inside brush
    front	342 (24.62mm/.969")     305 (25.52mm/1.005")
    back	339 (24.8mm/.9765")	3035 (25.6mm/1.008")
    
    Note heights are slightly higher on outside brush than previously.
    Jeff recommends these heights for optimum cleaning. Please do not 
    change these heights without approval from process engineering.
    
    Jeff also wrote a new cleaning recipe with faster brush rpm (called
    stras_clean_best).
    
47.12SOFTWARE UPGRADESUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0215
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"   12 lines   7-OCT-1994 15:58
                             -< Passdown for 10/7 >-
--------------------------------------------------------------------------------
    
      Ron came in and completed yet another software
    upgrade to this machine.  It was designed to 
    eliminate some small bugs that were irrevalent
    to us but none the less, they're gone.  The
    new rev. # is listed in the book.
    
      Ron also replaced the (4) spin station fingers
    to a bigger style with better wafer handling.
    Wafers were run through without any problems.
     
    BJ
47.13ROLLER ARMS NOT CLOSINGSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0211
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"    8 lines  13-OCT-1994 16:10
                            -< Passdown for 10/13 >-
--------------------------------------------------------------------------------
    
    
       The situtation with the wafer sliding past the 
    outter brush came back.  I increased the air pressure 
    on the rollers to try and stop this from happening.
    It has been fine all day so please keep an eye on it.
    
    BJ
47.14SCRUBBER WON'T GO INTO STARTSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0213
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"   10 lines  19-OCT-1994 16:55
                            -< Passdown for 10/19 >-
--------------------------------------------------------------------------------
    
    
    
        The scrubber wouldn't go into START.  I went over
    to check on the OCP and sure enough, the machine was in 
    an alarm condition.  Why? I could not find out why.  Mike
    from Ashland said everything was ok downstairs.  I cycled 
    the power to the OCP and everything was fine after that.
    
                                                     BJ
47.15HEAT LAMP NOT COMING ONSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0312
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"    9 lines   2-NOV-1994 16:23
                             -< Passdown for 11/2 >-
--------------------------------------------------------------------------------
    
    
       The lamp wasn't coming during its cycle.  The lamp
    was in the socket but wasn't making good contact.  I
    believe it was due to vibration in the room.  I will
    start to look into getting some kind of a clip to hold
    it into place.
    
                                              BJ
47.16SPIN STATION WAFER SENSORSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0312
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"    9 lines  21-NOV-1994 09:33
                            -< Passdown for 11/21 >-
--------------------------------------------------------------------------------
    
         The wafer was going into the SPIN station but
    the wafer sensor wasn't detecting the wafer coming 
    into the station.  I had to adjust the wafer sensor
    amplifier to dectect the wafer.  It just needed to 
    be tweaked.  I ran (6) wafers without any problems.
    
    
    BJ
47.17HEAT LAMP WOULDN'T COME ONSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0410
PATE::GEE                                             7 lines   5-DEC-1994 22:42
                              -< Heat Lamp Bulb >-
--------------------------------------------------------------------------------
    The heat lamp didn't turn on. The bulb keeps vibrating out of the
    spring clips, so I adjusted the spring clip bracket assembly in
    slightly to put more pressure on the bulb.
    
    Bob
    
     
47.18SPIN STATION HEADSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0411
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"    8 lines   9-DEC-1994 11:49
                             -< Passdown for 12/9 >-
--------------------------------------------------------------------------------
    
        The four screws on the cap that holds the
    fingers are suppose to be snug, not tight.
    We loosened the screws and the fingers slide
    nice now.  We bent the tabs in the lamp socket 
    so the lamp will stay tight.
    
                                       BJ
47.19A2 CHARACTERIZATION SHEET AS OF 12/12/94SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:05204
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"  201 lines  12-DEC-1994 03:03
                 -< A2 CHARACTERIZATION SHEET AS OF 12/12/94 >-
--------------------------------------------------------------------------------
   
                              CMP_SCRUB.A2
                        ONTRAK DSS-200 SCRUBBER
               MACHINE MAINTENANCE CHARACTERIZATION SHEET

LOAD AREA
==========

                                             ____
ELEVATOR: PHYSICAL POSITION                  ||_| <-- (1)
          (measurement taken from (1)top     ||    ^
          of the elevator mounting           ||    ^
          bracket to the (2)bottom of the    || | <-- (2)
          motor mount casing.                ||||_____ <--elevator base
           
 

                                UPPER POSITION (SLOT #1)  -- 1 7/8"  

                                HOME POSITION -- 1 3/16"

                                LOWER POSITION (SLOT #25) -- 7 25/32" 


MACHINE OFFSET --  (found in ROBOT OFFSETS menu)
                   
                                 LOAD STATION --  1250
  
              

SPRAY MAINIFOLD --  
(bar and nozzle position)      verify alignment of scribed marks

============================================================================

SCRUB MODULES
=============


MICROMETER SETTINGS:  verify ENGINEERING set point

                                 INSIDE BRUSH STATION | OUTSIDE BRUSH STATION
                                    SETTING/ACTUAL    |     SETTING/ACTUAL
                                 --------------------------------------------
                                                       |
WAFER ROTATION (TACH READING IN RPM'S)  40    /  44    |   50      /    45
                                                       |
BRUSH ROTATION (TACH READING IN RPM'S) 118    /  117   |   118     /    114


SPRAY MAINIFOLD --  
(bar and nozzle position)      verify alignment of scribed marks

==============================================================================

SPIN STAION
===========
                                            SETTING/ACTUAL 
                                            --------------
                        
WAFER ROTATION (TACH READING IN RPM'S)        1800 / 1815     

                                               500 / 624


CUP HEIGHT -- (measurement taken         UPPER POSITION -- 7 3/8"        
              from very top of any     
              finger to bottom of        MIDDLE POSITION -- 6 13/16"
              tub)
                                         LOWER POSITION -- 4 5/16"


GRIPPER FINGERS -- (measurement taken          FINGERS OPEN -- 8 19/32"
                   from center of left 
                   rear finger to center       FINGERS CLOSED -- 7 19/32"
                   of left front finger)


WAFER TEMPERATURE --  AS WAFER ENTERS SPIN STATION  -- 75 F

                      AS LAMP IS ON  -- 80 F

                      AFTER LAMP SHUTS OFF  -- 75 F


EXHAUST VELOCITY (in inches) -- 1.3


============================================================================

ROBOT
=====

OFFSETS --     WRIST AT SPIN STATION    --  64

               WRIST AT RECEIVE         --  304

               SPIN XPORT DELAY         --  42

               ROTATION AT SPIN STATION --  862

               ROTATION AT RECEIVE      --  1837

               ROBOT FINGERS            --  380


ROBOT FINGERS -- (measurement taken         OPEN -- 9 5/8" 
                 from O.D. to O.D.
                 from finger to finger      CLOSED -- 8 1/4"


==========================================================================

UNLOAD MODULE
=============


CASSETTE AT POSTION SLOT #1-- (insert new cassette
                              and let machine go to
                              position and slot #1.     --  16 9/16"
                              measure from the outer
                              most part of the back
                              panel the middle of slot
                              #1)
                               

KNIFE EDGE -- (measure from very         UPPER POSITION -- 10 1/4"
              top of knife to the
              top of panel that the      LOWER POSITION -- 7/8"
              edge goes into)



MACHINE OFFSET --  (found in ROBOT OFFSETS menu) 
                   
                                UNLOAD STATION --  442

                                BLADE LIFTER --    11650

=============================================================================

ELECTRONICS
===========
                                                  IDLE     |   RUNNING
                                                ========================

POWER SUPPLY VOLTAGE --               
                        +5 VOLT POWER SUPPLY       5.18    |   5.18
                                                           |
                        +12 VOLT POWER SUPPLY      13.1    |   13.1
                                                           |
                        -12 VOLT POWER SUPPLY     -13.8    |   -13.8  
                                                           |
                        +24 VOLT POWER SUPPLY      24.1    |   24.1
                                                           |
                        +48 VOLT POWER SUPPLY      48.2    |   48.2


SERVO DRIVE POWER SUPPLY-  
                       
                        +40 VOLT PICK-OFF          41      |   41
                                                       
                        +5 VOLT PICK-OFF           5.1     |   5.1


============================================================================

FLOW SETTINGS
=============

                                     SEND --  5

                  AMMONIA OUTSIDE STATION --  3

                      OUTSIDE BRUSH UPPER --  10

                      OUTSIDE BRUSH LOWER --  7

                   AMMONIA INSIDE STATION --  3

                       INSIDE BRUSH UPPER --  9

                       INSIDE BRUSH LOWER --  7

                       SPIN STATION UPPER --  3

                       SPIN STATION LOWER --  3

                        MEGASONIC NOZZLE  --  9

===========================================================================

PRESSURE SETTING                                50  PSI
================


SIGNET FLOW SETTING
===================                         LOW --  40

                                            HIGH -- 0

47.20LIFER MISSING WAFERS AND DOUBLE SLOTINGSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0620
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"   17 lines   3-JAN-1995 07:05
                             -< Passdown for 1/3 >-
--------------------------------------------------------------------------------
    
       I notice the lifter at the receive station missed a 
    
    wafer when looking for an open slot.  This would of caused
    
    a crash if it had gone undetected.  Sometimes it may be
    
    a bad slot in the cassette.  Adjusted the RECEIVE positioning
    
    variable to be more centered on the blade lifter.  It was at
    
    442 and I moved it to 430.  I let the machine do a wafer slot
    
    check on an entire wafer and all of the wafers were handled 
    
    fine.            
                                                   BJ
47.21double slotting wafers in the receive cassetteSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0615
PATE::GEE                                            12 lines   9-FEB-1995 23:34
                       -< receive station, weekly p.m. >-
--------------------------------------------------------------------------------
	Problem: System was double slotting wafers in the receive
                 cassettes,number 1 slot.

	Action: 
               * Inspected the receive cassette drive chain, no problems found.

	       * Rebooted system and ran 4, 6 wafer runs with no problems.

	       * Completed the weekly p.m.

						John & Bob
                           
47.22SOFTWARE UPGRADESUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0720
SUBPAC::LANDRY                                       17 lines  13-FEB-1995 17:17
                            -< Ron to the rescue! >-
--------------------------------------------------------------------------------
    Ron Simmons came in today to perform a couple of upgrades.  New
    software was installed.  It will prohibit the loss of data when the
    system gets shut off.  He also installed a booty in the spin station 
    to help prohibit any water from migrating into the lower cabinet.
    
    He showed up with impeckable timing today.  The system had just gone
    down for a lack of water to the upper brushes.  Seeing I was tied up
    with the Strasbaugh, he said he'd repair the problem.  A solenoid was
    replaced to fix this.  Thanks Ron!
    
    							Doc
    
    ps.  On both Scrubbers (A1 & A2), the new software installation wiped
         out all login / passwords.  If you had this capability, you will
         need to re-do them.
    
    
47.23coiled electrical cord/send station elevator SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0715
SUBPAC::LANDRY                                       12 lines  22-FEB-1995 16:36
                      -< Same cord problem as system A1 >-
--------------------------------------------------------------------------------
    The coiled electrical cord for the send station elevator has been
    chaffed conciderably.  It was beginning to create particles on the 
    outside of the plexiglass box (no contamination within the box where
    the wafers are loaded).  As a temporary provision to stop this from
    happening, a strand of electrical tape was applied.
    
    OnTrak will be notified of the problem.  We'll get the cords replaced
    as well as possible changing the "style" of hardware that protrudes
    from the plexiglass box.  It currently has hex-head hardware where a
    button head style would probably work out better.  Stay tuned!
    
    						Doc
47.24Robot fingers not repeating SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0818
SUBPAC::BJUBINVILLE "Brooks, Tritt and Diffie"       15 lines  14-MAR-1995 09:07
                        -< Robot fingers not repeating >-
--------------------------------------------------------------------------------
    
    
          It started out as an easy semi-annual PM and it just got 
      worse.  The PM was completed but upon checking the finger 
      operation on the robot when I noticed the fingers not repeating
      and barely holding onto the wafer.  Curiosity lead to removing the
      covers around the robot and the linkage that the OPEN and CLOSED
      flags are was all loose.  After the clamp was tightened down, the
      finger got a little better but still not right.  After speaking
      with the field servive rep., he agreed on the steps that were taken
      as well as trying some other things.  Doc took the bull by the horns
      and is having a rough ride.  Ron will be here in the morning if
      we can't solve this issue.
                                        BJ and the Doctor
    
47.25The robot with inconsistant fingers SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0840
SUBPAC::LANDRY                                       37 lines  14-MAR-1995 17:06
                    -< The robot with inconsistant fingers >-
--------------------------------------------------------------------------------
    		The Robot with the Inconsistant Fingers
    
    			     Chapter II
    
    We left off in chapter I with BJ returning to the ranch after a
    long, hard evening and Doctor grasping the reins on the bull-
    headed robot.
    
    The robots fingers were not consistant in operation.  They would grab
    the wafer correctly at times, sometimes not.  The sensor would recog-
    nize the wafer, sometimes not.  The flag that interrupts the sensor
    would end up in different locations, only rarely repeating itself.
    Found that the fingers open photo sensor was repeatable every time.
    The fingers closed photo sensor, on the other hand, was incredibaly
    inconsistant.  Cleaned the closed photo sensor off to assure proper
    opticle performance, no help.  Checked the resistor / pc board and
    again no help.  
    
    Knowing that the open fingers sensor was repeatable, began making
    mechanical adjustments to the close fingers sensor's flag.  We began
    to get some different results.  It appeared that these adjustments just
    may do the trick as things were improving.  As we slightly moved the
    flag again and retried the robots performance, it picked up the wafer
    fine.  Then as the wafer was down-loaded into the cassette, the system
    timed out in search of the fingers open sensor.  The flag had flown
    right on by the sensor.  Back to square one.
    
    The motor drives the fingers using step counts for location.  The photo
    sensors are only used to say that the task has been completed and the
    next function may now take place.  
    
    There are several components that may be causing this, PIA or Int'l
    stepper board, photo switch flakey, connections/cables.  It could even
    need the robot functions re-taught.  The rep is due in Wednesday
    morning to look at the problem (warranty work).
    
    							Doc
47.26robot finger errorSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0812
SUBPAC::BJUBINVILLE "Brooks, Tritt and Diffie"        9 lines  15-MAR-1995 05:59
                            -< robot finger error >-
--------------------------------------------------------------------------------
    
        After continuing to make adjustments to the robot, it finally
    started working.  It was just a matter of being really touchy. 
    Yesterday, Doc and I noticed some software problems that need
    to be addressed by Ron Simmons.  Ron is suppose to be here this 
    morning.  There is no work for the machine so I left it down.  
    If it is needed, it can be run.  The software problem won't
    effect the wafers.
                                             BJ
47.27brushes NOT STAYING ON IN IDLESUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:0922
SUBPAC::LANDRY                                       19 lines  15-MAR-1995 16:16
                        -< To rotate or not to rotate! >-
--------------------------------------------------------------------------------
    System status..........
    
    BJ did a nice job of bringing in the robot gripper sensors to where
    all looks fine.  Ran all day without a hitch.
    
    At the start of the shift, we couldn't get the brushes to rotate in
    it's normal idle state.  By using the maintenance screen I/O func-
    tions, we manually turned the brushes on.  This would suffice until
    the system would run it's cassette of wafers and end.  Upon removing
    the cassette, the brushes would shut down (as should) but after putting
    the system in start, the brushes were still (no rotation).  Back into
    maintenance and manually start them.  Well, the next time they ran the
    system (and everytime after that) the brushes would rotate in every
    state they normally would.  So, I guess we'll keep an eye out for this
    phenomina to return!
    
    Bottom line, alls well for know.
    
    							1st shifters
47.28robot finger errorSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1014
SUBPAC::LANDRY                                       11 lines  16-MAR-1995 10:47
              -< The case of the robot fingers has been solved! >-
--------------------------------------------------------------------------------
    Today..........
    
    Ron Simmons came in today to check out the problem we experienced this
    week with the robot.  Although the system was available for manufactur-
    ing, the robot performance wasn't back to it's normal functionality.
    Therefore, he re-taught the gripper function positions and re-setup
    the photo sensor flags as well.  The "bounce" that was left in the 
    gripper movement has vanished and the robot is humming along fine.
    
    							Doc
    
47.29MACHINE OFF-SETS AS OF 3/21/95SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1026
SUBPAC::BJUBINVILLE "Brooks, Tritt and Diffie"       23 lines  21-MAR-1995 06:06
                      -< MACHINE OFF-SETS AS OF 3/21/95 >-
--------------------------------------------------------------------------------
    
    
CMP-SCRUB.A2 MACHINE OFF-SETS AS OF 3/21
    
    
               LOAD STATION               --  1250
   
               WRIST AT SPIN STATION      --  64
       
               WRIST AT RECEIVE           --  304

               SPIN XPORT DELAY           --  42

               ROTATION AT SPIN STATION   --  864

               ROTATION AT RECEIVE        --  1837

               ROBOT FINGERS              --  270

               UNLOAD STATION             --  430

               BLADE LIFTER               --  11650
                                             
47.30Load Station elevator wiring harness replacedSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1119
SUBPAC::LANDRY                                       16 lines   7-APR-1995 14:16
                          -< wiring harness replaced >-
--------------------------------------------------------------------------------
    Ron Simmons, OnTrak Rep Extraodinaire, came in today to replace the
    wiring for the Load Staion elevator power.  The original wiring was
    mounted such that the wiring would rub against the hardware protruding
    from the load station tub.  The new wiring has been installed such that
    it will no longer rub in the previous manner.  The socket head screws
    will be replaced in the near future to button head screws, hence
    preventing this chaffing to occur again.
    
    We also checked out the possible leak on the feedthroughs in the brush
    stations.  Water was noticed to have formed there on one of his
    previous visits.  The last two times he was here, they have been dry,
    as should be.  Original sighting could have possible been attributed to
    wet transfer boats, as they are some times temporarily placed there.
    We'll monitor for future sightings.
    
    						Doc
47.31Send Station Rollers/Bad Bearings SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1117
ASDG::POIRIER                                        14 lines  10-APR-1995 19:23
                     -< Send Station Rollers/Bad Bearings >-
--------------------------------------------------------------------------------
    
    * Send Station Track Problem:
    
      The send Station Track was moving much slower that normal on 
      first shift today. The system was alarming for "Time out" errors
      at the Outside Brush Station. Bobby disassembled the Send Station
      roller assembly and discovered high friction in the bearings. The
      motor could just barely over-power the bearings. Bobby is planning
      to replace the bearings on 2nd shift tonight.
    
    
                                                            John & Bob.
      
     
47.32send station belt bearings SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1110
PATE::GEE                                             7 lines  10-APR-1995 23:10
                        -< send station belt bearings >-
--------------------------------------------------------------------------------
	Problem: The wafers wouldn't go to the outside brush station.

	Fix: Replaced two bearings on the send station belt drive. They
             were all rusted and not functioning. Ran 50 wafers with no 
             problems, ran a particle check 2 adders. 

						Bob
47.33Roller arm was loose SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1110
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               7 lines  12-APR-1995 06:06
                           -< Roller arm was loose >-
--------------------------------------------------------------------------------
    
    
        The front roller in the outside brush station was opening
      during the scrub cycle just from the force of the wafer.  
      The arm that clamps down on the shaft had all of its screws
      loose.
                                                   BJ
47.34broken roller armSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1215
SUBPAC::LANDRY                                       12 lines  12-APR-1995 17:54
                     -< No cast will fix that broken arm! >-
--------------------------------------------------------------------------------
    The system is down for parts.  One of the roller arms in the outside
    brush station broke today.  Without the use of the arm, the wafer would
    blow right on by the brushes and not get scrubbed at this station.  
    Well, there is no way we have an item of that nature stocked here.  So,
    the rep was contacted.  He believes he has the arm here on the east
    coast.  He will be in tomorrow morning to replace the defective one.
    If he has to get the part from the west coast, we're probably looking
    at another day to get it shipped.  The rep is researching the situation
    as I type.
    
    						Doc
    
47.35repaired broken roller armSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:129
SUBPAC::LANDRY                                        6 lines  13-APR-1995 19:31
                         -< Spare Parts, the sequel! >-
--------------------------------------------------------------------------------
    Ron Simmons arrived today to repair the broken arm on the roller
    assembly.  Upon attempting the task, it was noted that additional parts
    would be needed (a second arm).  He is having them ship them out
    priority from California.  We should receive them in the morning.
    
    						Doc
47.36broken roller armSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1210
PATE::GEE                                             7 lines  20-APR-1995 00:01
                                   -< Arm  >-
--------------------------------------------------------------------------------
	Problem: Wafer would not exit the inside brush station. Found
                 the front brush arm broken at the shaft bolt holes.

	Fix: Found a replacement in the Speedfam IO cabinet. Replaced,
             realigned and preformed a verification particle test it
             all set. 
						Bob
47.37replaced cracked roller arm SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:137
SUBPAC::TOOKER                                        4 lines   2-MAY-1995 15:44
                        -< replaced cracked roller arm >-
--------------------------------------------------------------------------------
    Inside brush station back roller arm cracked.  Installed a new arm
    and ran wafers to verify operation.
    
    Jim T.
47.38DI water gun was installed SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1315
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              12 lines   9-MAY-1995 05:45
               -< Weekly PM completed, DI gun installed for ENG >-
--------------------------------------------------------------------------------
    
    
       Weekly PM completed....
     
       Note:  A DI water gun was installed on this machine.  It
              picks off the right main 3/8" line and is shut off
              by a needle valve.  To use the gun, simply open the
              valve.  When you are done, shut the valve off and drain
              any DI water from the line.  Then, return the gun to
              where you found it.
                                                     BJ
               
47.39Megasonics sweep arm sensor SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1312
PATE::GEE                                             9 lines  10-MAY-1995 23:23
                        -< Megasonics sweep arm sensor >-
--------------------------------------------------------------------------------
	Problem: The Megasonic arm in the spin station would not
	         move to home position.

	Fix: Found the set screw that holds the sweep arm sensor flag
	     had loosed up. We retightened the set screw and ran 20
	     wafers through with no problems.
					

			Second Shift
47.40Outer Brush motor problemSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1412
FABSIX::R_POIRIER                                     9 lines  15-JUN-1995 23:42
                             -< Outer Brush motor >-
--------------------------------------------------------------------------------

There was a problem with the Outside top brush motor. It kept 
stopping when in idle mode or it would stop if any torque was 
applied to the motor, and would not start again by itself. We
replaced the motor and it has been running ok since.


					2nd Shift

47.41new motor spin speed calibration SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1422
SUBPAC::LANDRY                                       19 lines  16-JUN-1995 18:04
                     -< new motor spin speed calibration >-
--------------------------------------------------------------------------------
    This morning...........
    
    	I took a few minutes to see exactly what the outside brush station
    top brush was spinning at since the new motor installation.  It was 
    within spec, but I attempted to bring it in even tighter.  I found that
    the calibration procedure (which is located in the OnTrak manual as
    well as the Digital Maintenance spec) was rather hard to follow and
    execute.  Therefore, I contacted Ron Simmons for some guidance.  He
    clearified a few "hazy" spots for me.  I will edit the maintenance
    spec soon to better explain some of the circumstances.  Before I go to
    far, let me say that the procedure is correct in both the spec and the
    manual.  The planned editing will make it simplified.
    
    Anyway, I used the tachometer to vefify that we were in spec.  At a
    setting od 125 RPM (cleaning step speed!) the external tachometer was
    reading 122.4 RPM, well within the 10% variance allowed.
    
    						Doc
    
47.42OB sta brush change, new micrometer settings!SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1430
SUBPAC::LANDRY                                       27 lines  23-JUN-1995 16:50
               -< OB sta brush change, new micrometer settings! >-
--------------------------------------------------------------------------------
    
    	The infamous leak was repaired by the 3rd shifters.  It's holding
    the water back real well.  No runs, no drips, no errors.....one part
    left over in the fab.  I found a small o-ring on the system.  Not sure
    if this was a changed out one, one left out or what.  But hey, there
    are no leaks, so I "ass/u/me" that all is well???
    
    Moved on to the new brush installation.  This is usually an arduous
    task to say the least.  I don't know what new technique I used, but it
    seemed to go fairly smoothly.  Got the nubs to stay straight, evenly
    distributed too!  Re-installed said brushes.  Ran a full cassette of 
    "break-in" wafers.  Earl was ever_so_kind to perform the particle
    checks/MQC.  Results were positive.  The systems back on line.
    
    Only the outside station had the brushes changed out, as these were the
    only ones to get contaminated.  All 3 polishers were running, hence the
    return of the scrubber to manufacturing.  As soon as time permits, the
    inside brushes can get changed.  We need to keep the brush life between
    stations consistant with each other.  Once those brushes are changed,
    we'll need to log down the wafer count in the NOTES file, then reset
    the count on the system.
    
    One last thing..........the new micrometer settings for the outside
    brush are:  Front = 3.05 + Rear = 3.10.  This info has also been
    entered in the log book on the system.
    
    							Doc
47.43New Brush Height Micrometer Settings SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1424
SUBPAC::LANDRY                                       21 lines  24-JUN-1995 15:51
                   -< New Brush Height Micrometer Settings >-
--------------------------------------------------------------------------------
    Between Friday morning and Saturday morning, all four brushes were
    changed out on the system.  The brush heights were adjusted.  The new
    micrometer settings are:
    
    		Outside Station			Inside Station
               -----------------               ----------------
    
    		Front = 305			Front = 310
    		Rear  = 310			Rear  = 310
    
    Please use these new settings when checking the brush heights during
    the PM's.  he numbers can also be found in the logbbok on the system.
    
    Break-in wafers were run by third shift.  MQC's were run on first
    shift.  Minor adjustments needed to be made to the micrometers after
    the break-in wafers were run.  Also adjusted the roller arms.  The
    wafer stop position (which the rollers control) were out of adjustment,
    letting the wafer over travel a bit.  
    
    						Doc
     
47.44Recieve Station Door Sensor SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1510
ASDG::POIRIER                                         7 lines  25-JUL-1995 19:34
                        -< Recieve Station Door Sensor >-
--------------------------------------------------------------------------------
    
    * Recieve Station Door Sensor:
    
      Since the sensor was undamaged electrically, its' plastic housing was
      bonded back together with Loctite. The door hinges were then
      readjusted to compensate for the sag in the door. The sensor
      broke due to interference with the door. 
47.45broken or lost wafer errorSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1518
SUBPAC::LANDRY                                       15 lines  29-JUL-1995 18:19
                     -< broken or lost wafer error....... >-
--------------------------------------------------------------------------------
    While checking out the "alarms" log on the system today, I noticed
    several errors for "broken or lost wafer" in the spin station.  They
    are all fantom errors as the system never stops running.  While
    watching the system outputs page, I noticed that occasionally, the
    spin fingers closed LED would flash on then off and remain off.  To
    familiarize you with this sensor, the fingers closed LED should NOT
    light up, as this signifies that the fingers closed to their mechanical
    end, and not the position that the wafer is grabbed.  Therefore, when
    the "fingers open LED is lit", this is a true sensor reading.  Once the
    fingers close around a wafer, this LED should go off and the Closed LED
    should also remain off. So, I mechanically adjusted the microswitch for
    this fix.  We'll have to monitor it to see if it is adjusterd enough.
    
    							Doc
    
47.46PROCESS ENGINEERING CHANGESSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1530
SUBPAC::LANDRY                                       27 lines   4-AUG-1995 19:20
                   -< major adjustments made / per Pro Eng >-
--------------------------------------------------------------------------------
      Today, Frank Krupa and scrubber extrodinaire Pat McPherson (from fab 4)
performed some expirements on the system.  They did the bulk of the job and
I assisted them with equipment adjustments when necessary.  The objective was 
to improve the megasonics performance and reduce particles.  Many different 
parameters were tried, as well as equipment adjustments.  Some of the physical 
adjustments showed extremely positive results with particle removement.  They 
include the raising of the megasonics arm, changing of the water flow rate at 
the arm, increasing of the megasonics generator.  Also, spin speeds helped 
things dramatically.  Due to operating specifications, some of the changes 
could not be made perminent at this time.  But, what items were not affected 
by the spec have been made perminent.  They are the Megasonics arm height and 
generator setting and also the flow rate from the megasonics nozzle.  The 
height has been changed to 3/8" clearance over the wafer and the generator 
has been turned up to it's max setting.  The flow rate has been reduced to 
500ML (or 5 on the flow meter).

      Other changes to be made in the near future as the spec is DCO'd.


      On another note, the "wafer lost or broken" alarm has been repaired.
The actual switch that was out was the reed switch on the bimba shaft.  This
was detected by our ace Equipment Engineer, Jim Tooker.  With his guidance,
we got the switch adjusted right (I hope  8*> ).

      						Doc                      


47.47BRUSH CHANGES AND ALIGNMENTSSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1638
LUDWIG::KRUPA                                        35 lines  23-AUG-1995 17:59
                       -< Ontrak Training/brush set-up >-
--------------------------------------------------------------------------------
WED - 08/23/95
Ron Simmons from Ontrak worked on CMP-SCRUB.A2 today.  John Pourier, Doc,
and Brendan Finn, and myself were trained by Ron to perform the following:
	1 brush change procedure: NOTE: NEED A BUCKET IN CMP ROOM TO SUBMERSE
					THE BRUSH CORE IN TO MAKE INSTALLATION
					EASIER.
	2 brush parallelism: WERE WAY OUT ON BOTH TOOLS
	3 brush height set-up: BOTH BRUSHES WERE SET TOO HIGH
	4 alignment of wafer transfer in spin station and unload station

NOTE: Procedures 1-3 were also performed on CMP-SCRUB.A1 today.  Procedure 4 
	was not due to time and system availability, but, needs to be done.

In addition Ron performed a "once-over" on the tool and identified the
following issues; some of which were corrected:
	- No NH4OH was flowing to SCRUB.A1&A2.  All flow regulators were closed.
	  even when they were opened, there was no supply and NO ALARMS!
	  Doc called Ashland chemical.  Shortly after the call we heard the low
	  flow alarm go off.  Soon after the NH4OH began to flow.  Plan to
	  look into this further tomorrow to determine how long this was this
	  way and why we had no alarms.
	- The exhaust for the spin station and brush stations are not set-up
	  correctly(their required exhaust tags are actually reversed).
  	  Brush station should be ~5 scfm.  Spin station should be ~50scfm.
	  Actual was ~16 scfm for both.
	- The water in the spin station was turning on prior to the carriage
	  dropping into low position.  This caused water to splash off the
	  wafer and out of the bucket.  The low position sensor was found to
	  be bad(~1.2mV) and was replaced with a new one(0.26mV).
	- The wafer appears to "wobble" when spinning in the station. 
	  Alignment of the mechanism did not appear to help this any.  Ron
	  will be looking into causes, but, we should be ok for now.
	- The top rinse was not centered on the wafer and this was corrected.

-Frank, Ron, Doc, John, Brendan
47.48BRUSHES ALIGNMENTS, CHANGES AND EXHAUSTSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1637
SUBPAC::LANDRY                                       34 lines  23-AUG-1995 19:20
                   -< Brushes changed out.  Sensor changed. >-
--------------------------------------------------------------------------------

WED - 08/23/95
Ron Simmons from Ontrak worked on CMP-SCRUB.A2 today.  John Poirier, Doc,
and Brendan Finn, and myself were trained by Ron to perform the following:
	1 brush change procedure: NOTE: NEED A BUCKET IN CMP ROOM TO SUBMERSE
					THE BRUSH CORE IN TO MAKE INSTALLATION
					EASIER.
	2 brush parallelism: WERE OUT 
	3 brush height set-up: BOTH BRUSHES WERE SET TOO HIGH
	4 alignment of wafer transfer in spin station and unload station


In addition Ron performed a "once-over" on the tool and identified the
following issues; some of which were corrected:
	- No NH4OH was flowing to SCRUB.A2.  All flow regulators were closed.
	  even when they were opened, there was no supply and NO ALARMS!
	  Doc called Ashland chemical.  Shortly after the call we heard the low
	  flow alarm go off.  Soon after the NH4OH began to flow.  Plan to
	  look into this further tomorrow to determine how long this was this
	  way and why we had no alarms.
	- The exhaust for the spin station and brush stations are not set-up
	  correctly(their required exhaust tags are actually reversed).
  	  Brush station should be ~5 scfm.  Spin station should be ~50scfm.
	  Actual was ~16 scfm for both.
	- The water in the spin station was turning on prior to the carriage
	  dropping into low position.  This caused water to splash off the
	  wafer and out of the bucket.  The low position sensor was found to
	  be bad(~1.2mV) and was replaced with a new one(0.26mV).
	- The wafer appears to "wobble" when spinning in the station. 
	  Alignment of the mechanism did not appear to help this any.  Ron
	  will be looking into causes, but, we should be ok for now.
	- The top rinse was not centered on the wafer and this was corrected.

					Frank, Ron, Doc, John, Brendan
47.49sensor replaced/motor speeds changedSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1712
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               9 lines  23-SEP-1995 09:01
                   -< sensor replaced/motor speeds changed >-
--------------------------------------------------------------------------------
    
        The home sensor for the SEND station blew out and was
      replaced.
    
        The OUTSIDE brush and roller speeds were slightly out of spec
      and kept tripping the alarm.  The speeds were set to run at the
      program speeds and the % error was brought to 0.
    
                                                        BJ
47.50Send Station Top limit sensorSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1710
PATE::GEE                                             7 lines  26-SEP-1995 07:16
                       -< Send Station Top limit sensor >-
--------------------------------------------------------------------------------
	The send cassette top limit switch wasn't working.
 
	Found that the sensor is cracked. I repaired it for now,
	but didn't have time to replace it. I pulled one from stock,
	it's onto of the tool box.
	
					B Shift
47.51Input Cassette SensorSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1715
FABSIX::B_FINN                                       12 lines  26-SEP-1995 18:40
                           -< Input Cassette Sensor >-
--------------------------------------------------------------------------------


	Replaced input cassette top stop switch.  (Please refer to B shifts 
previous note on this tool).  Adjusted sensor's flag so it will not strike 
the new sensor. Ran several times to test, also ran half a boat of wafers, aok.
  
The old switch which is cracked, but functional (in a crisis) was tagged and 
placed in the tool box.

					A Shift

			
47.52Outside station Roller speedSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1742
PATE::GEE                                            39 lines   3-OCT-1995 07:25
                       -< Outside station Roller speed >-
--------------------------------------------------------------------------------
	When we came in tonight the system was down for an 
	"outside brush-roller motor speed failure". The speed
	was set at 50 rpm but the display only read 33 rpm.

	Talked to A shift and they informed use that the procedure
	in the manual didn't work. We called the field service rep.
	to see if we could get a FAX of a procedure that works.
	
	He told use to try and adjust pods R31 and R13 and see if
	we could bring the speed into spec. It didn't do anything.
 	Pod R31 is located on the main motor controller board and 
	controls the speed in all areas of the system, we put the 
	Pod back to its original position. Called a second time and 
	never received a response.
  
	Check the 5v,12v,24v and 48volt power supplies and they were
	all good. Checked the inside brush station speeds and they 
	checked good also.
	
	We decided to try the procedure in the manual on calibrating 
	the Dual DC motor controller board for motor drive #4, 
	this controls the outside brush-roller motor. When this
	procedure is completed the set speed of 50 rpm should
	be displayed on the screen and by adjusting pod R54
	you can adjust the roller speed to match the display and
	setpoint. 
	At a setpoint of 50 rpm we displayed 0 rpm and at the roller
	we read 20 rpm.

	We replaced the board hoping that one of the pods may have been
	stuck, no luck we had the same results.
	
	By adjusting pods R54 and R13 we could change the speed at the 
	roller and on the display screen, but never by the same amount.
	
	Need to contact field service in the am for a procedure on 
	calibrating the speeds.
	
					B Shift
47.53ROLLER MOTOR tach problems SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1827
SUBPAC::LANDRY                                       24 lines   3-OCT-1995 19:35
                               -< tach problems >-
--------------------------------------------------------------------------------
    The ride continues to be rough.......
    
    *  With the assistance of Ron Simmons, the speed out of spec problem
       was found to be a bad tach on the roller motor.  The motor was
       replaced and board calibration for the motor was performed.
    
    *  The cause for the motor failure was due to a drip coming from the 
       outside station roller motor feedthrough, which was landing directly
       on the motor.  We RTV'd the leak and let it cure.
    
    *  Upon system verification run, the Spin Station motor was all out of
       whack, running away in fact.  This is where we are currently handing
       over the system.  Some info goes:
    
    	-  we input a spin speed of 600 RPM.  The tach readout on the screen
           showed 815 RPM.  The hand held tach read 768.  This is giving
           the system a out of spec tolerance.
    
    	-  Peter Michaels (California) is FAX'ing a procedure to correct
           this.  Also, Ron Simmons has been left a phone message about the
           problem.
    
    						A shifters
    
47.54Spin station / Send station SPEEDSSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1814
PATE::GEE                                            11 lines   4-OCT-1995 07:18
                        -< Spin station / Send station >-
--------------------------------------------------------------------------------
	* The spin station spin speed was out of spec.
	
		Calibrated the Spin station spin speed. We used
		the procedure in the manual.

	* Wafers not arriving at the outside brush station.

		Found the wafer sensor in the send station loose, 
		tightened the sensor and adjusted the off-set variable.
		
					B shift & Brian
47.55transport rebuilding/ CARRIAGESUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1816
SUBPAC::LANDRY                                       13 lines   4-OCT-1995 18:18
                      -< pm, transport rebuilding....... >-
--------------------------------------------------------------------------------
    *  Performed the Weekly PM; check sheet looged into the proper note.
    
    *  Rebuilt a defective transport pulley in the outside brush station.
       All parts are stocked except for the pulley itself.   I will issue
       a SID sheet to stock these, as the one in the system shows some
       wear at the bearing entrance.  
    
    *  The leak in the outside brush station roller feedthrough tube that
       was repaired the day before, appears to have sealed nicely.  Still
       no evidence of a leak.
    
    						the A team
    
47.56Small D.I. Water Leak on Gage GuardSUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:1914
                    -< Small D.I. Water Leak on Gage Guard >-
--------------------------------------------------------------------------------
    
    * D.I. Leak:
    
      A small D.I. water leak has developed on the gage guard behind the
      system. It's on the pressure gage to the left as viewed from the rear
      of the tool. I attempted to order a replacement gage guard from the
      094 stockromm but the SGMR system was down. "D" shift, please replace
      the gage guard if time allows.
    
                                                               Thanks,
    
                                                                   John.
47.57gauge guards removed, amplifer issue SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:2014
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              11 lines  19-OCT-1995 06:31
                   -< gauge guards removed, amplifer issue >-
--------------------------------------------------------------------------------
    
    
         Removed both gauge gaurds on the DI feed.  Guess
      what...  We've got 40 psi!
    
          The megasonic arm didn't go to home.  It appears the 
      amplifier for the home sensor was not seated properly.  The
      arm move fine after that.
    
                                         BJ                                   
    
47.58Send Station Lock-up Problem SUBPAC::BJUBINVILLENFD Fire - RescueSat Dec 09 1995 00:2014
ASDG::POIRIER                                        11 lines  21-OCT-1995 20:01
                       -< Send Station Lock-up Problem >-
--------------------------------------------------------------------------------
    
    * Send Station lock-up Problem:
    
      The system errored numerous times this afternoon for "outside
      brush station timeout". The send station wasn't sending any wafers 
      out. We reset both the outside brush station and send station 
      numerous times only to get the error again. Brad and I finally
      EMO'd the system and rebooted. The system has performed numerous 
      runs without any further trouble.
    
                                                        John & Brad.
47.61arm not at homeFABSIX::B_FINNWed Dec 20 1995 18:389

        Megasonic's arm not homing.  Home position sensor not illuminated even 
when flag was met.  Found yellow sensor not seated properly in bracket. Also 
tightened arm to eliminate slop at top of arm as compared to Scrub.A1  Ran 
two boats of wafers aok.

                                  
						A SHIFT
47.62SYSTEM OFFSETS AS OF 12/21/95SUBPAC::BJUBINVILLENFD Fire - RescueThu Dec 21 1995 00:3620
    
               LOAD STATION               --  1660
   
               WRIST AT SPIN STATION      --  72
       
               WRIST AT RECEIVE           --  304

               SPIN XPORT DELAY           --  45

               ROTATION AT SPIN STATION   --  864

               ROTATION AT RECEIVE        --  1837

               ROBOT FINGERS              --  480

               UNLOAD STATION             --  430

               BLADE LIFTER               --  11600
                                             
47.63Send Elevator Sensor Adjustment / Timeout ErrorsASDG::POIRIERFri Dec 22 1995 17:5513
    
    * Send Elevator Sensor:
      --------------------
      The system errored numerous times today for "Timeout At Outside
      Brush Station". The Send elevator would lower the cassette until  
      the bottom most wafer contacted the belt. At this point nothing
      more would happen until the system errored for the timeout condition.
    
      The Send Elevator sensor was rotated counter clockwise by 1 turn.
      The sensor is now slightly higher/closer to the cassette. We've run
      approx 100 wafers thru the system since the sensor was adjusted
      without any timeout errors. Keep a watch on this condition as the 
      sensor may need a little more adjustment. 
47.64WAfer transport not workingSTRATA::RIDLONEliminate the obviousSat Dec 23 1995 06:545
    Wafer transport not working in send cassette. Sensor was seeing wafer
    but motor was not turning on. Motor worked fine in manual. Cycled power
    to reboot and ran some test wafers through with no problems.
    
    						MR
47.65Start-up for 12/29SUBPAC::BJUBINVILLENFD Fire - RescueFri Dec 29 1995 15:119
    
    
        The DI came on but is still out of spec so we didn't want
      to keep it to the tool so we shut off the DI feed and opened the
      by-pass.  Facilities still have some by-passes opened, that's
      why we don't have very high pressure.
    
     
                                                       Mitch,John,Leon,BJ
47.66CMP-SCRUB.A2 START-UP COMPLETESUBPAC::BJUBINVILLENFD Fire - RescueSun Dec 31 1995 14:4028
    
                CMP-SCRUB.A2 STARTUP CHECKSHEET
                -------------------------------

                                                       Shift       Date
                                                       -------     -------
  1.  Ensure that all necessary facilities are         B/C/D       12/31/95 
      available and ready to use.                      

  2.  Power up each system.                            B/C/D       12/31/95

  3.  Open incoming DI Water Valve to each             B/C/D       12/31/95
      system.                                          

  4.  Close UDR valve at each system.                  B/C/D       12/31/95

  5.  Replaced scrub brushes in outside                B/C/D       12/31/95
      brush station.  Re-installed brushes
      in the inside brush station.
    
  6.  Change RTC status to SCHEDULED DOWN and
      perform a Weekly PM on each system.              

  7.  Change RTC status to VERIFY and notify
      Operations that the systems are ready
      for MQC.                                       
                                                       
                                            
47.67ammoniaFABSIX::R_GEETue Jan 02 1996 07:493
	* Ammonia was down from 8 pm to 3am.

					B shift
47.68double loadedSUBPAC::LANDRYWed Jan 03 1996 17:468
    
    The system did something unnatural today.  It faulted out in the spin
    station.  Upon finding out why, I found two wafers in that station.
    No explanable reason why.  Removed them and tested system performance.
    Alls well.
    
    					Doc
    
47.69Arm replacedSUBPAC::BJUBINVILLENFD Fire - RescueWed Jan 10 1996 06:204
    
        Replaced rear top roller arm in the outside brush station.
    
                                      BJ
47.70brush hight adjustmentFABSIX::B_FINNFri Jan 19 1996 19:0611
  	   
              Found brush hights off. 

              ALL brush heights were adjusted new values are as follows;

	      Outside front = 375   Outside back = 355

	      Inside  front = 360   Inside  back = 335


						Frank/Brad/Brendan
47.71Brush height adjustmentSTRATA::RIDLONEliminate the obviousTue Jan 30 1996 00:024
    During the weekly pm found the inside back brush height off again.
    Adjusted to 335 and locked it in.
    
    					B shift
47.72 o-ring or drive belt?FABSIX::L_SHKOLNIKOVThu Feb 01 1996 07:1810
    * O-Ring in outside brush station
      --------------------
      The system errored numerous times today for "Timeout looking for 
      outside brush wafer sensor." A wafer went past the sender and stopped
      before outside brush station. Conveyer pulley wasn't spinning at the
      brush station. Found o-ring (202-113) was loose. Replaced o-ring.
      The machine worked fine after that.


                                           D shift
47.73need volts? oh here they are!SUBPAC::LANDRYTue Feb 06 1996 19:2410
Was asked to look at the system today for no water flow in the
Send Station.  Well, there was indeed water going to the station, 
but the 24 VDC solenoid wasn't energizing.  Checked for voltage
at the solenoid connector (J2).  24 Volts were there.  Just about
to order a new solenoid when something caught my eye.  J2 into J2!
I don't think so.  J2 should be plugged into P2.  After plugging
the proper sequence together, voila, water flow.

      					A Shift

47.74outside brush station lower brush bearingsFABSIX::R_GEESun Feb 11 1996 06:585
	* The outside brush station's lower brush was hesitating.

		Replaced all 4 bearing. 

					B shift
47.75System haltedSTRATA::FINNSun Feb 11 1996 19:1314
    
    
	System stopped dead in its tracks with a wafer at spin station
and fingers closed around wafer, Robot at receive position. Load cassette was 
not indexing and belts off. System would not retrieve wafer via wafer recovery 
screen. Check I/O for all station, everything seemed satisfied.  Tried to use 
I/O's to remove wafer, but system would not respond. Powered system down, upon 
boot up system prompted us to remove wafers via wafer recovery, this time it 
worked fine.  Ran 4 boats of wafers without error, returned to manufacturing.




						A shift 
47.76OCP failureSUBPAC::LANDRYTue Feb 13 1996 19:359
Today, the OCP for the Ammonium Hydroxide kicked off, sending both
scrubbers into a halt mode.  Checked the pressure gauge and it had
40 PSI going to the machine.  So why did they alarm.  Well, the 
fault displayed a "timed out" (or something wordy like that).  We
weren't able to reset the OCP's.  Gave Ashland a call to check out
there end of the deal.  The computer that controls the OCP hichup'd.
They need to down load a couple of files and we were back in business.

      					Doc
47.77motor replacedSUBPAC::LANDRYWed Feb 14 1996 19:3115
    Well, the system was aggravating the people in the room as the alarm
    horn kept sounding off for a "brush motor out of spec".  The upper
    brush in the Inside Brush station was the culprit.  High-tech'ly
    trouble shot the problem and found that the motor was bad.  It would
    run a programmed speed for awhile, then suddenly drop 20 RPM or so
    for a few seconds.  It was consistantly doing this.
    
    Therefore, the motor was replaced.  It ain't no fun trying to solder
    them little wires/caps when the motor is round and rolling on the
    table!  Once done, the motor was installed.  Of course, this means the
    board needed to be calibrated, which I did do.  Ran several exercise
    wafers to confirm the fix.  Okey Dokey!
    
    						Doc
    
47.78LeakSUBPAC::BJUBINVILLENFD Fire - RescueThu Feb 15 1996 06:316
    
        DI is leaking down the outside brush station at lower
      brush motor.   We will try and get the tool tomorrow
      morning if no one else gets a chance to get it.
             
                                     D shift
47.79Contact phone numbers for OnTrakSUBPAC::TOOKERFri Feb 16 1996 10:5027
Here is the latest list of service contacts for OnTrak Systems.  These con-
tacts can be used to get answers to technical questions, field service requests,
and parts.


    1)  Ron Simmons:  Office - 508-643-2772
                      Pager  - 508-504-7906 (this is your best bet)

        If no response in one hour, contact area supervisor in Fishkill, NY.


    2)  Pat Paino (area supervisor), Fishkill, NY:

                      Pager - 800-963-5770

        If no response in one hour, contact area supervisor in Austin, TX.


    3)  Stan Jaskinski (area supervisor in Austin, TX:

                      Office - 512-448-4494
                      Pager  - 800-905-9275

    4)  Brad Clements (regional manager), Austin, TX: (best bet for nights)

                      Office - 512-448-4494
                      Pager  - 800-796-7363  PIN 193-3639 (7 X 24 response)
47.80BRUSH MOTOR REPLACEDSUBPAC::BJUBINVILLENFD Fire - RescueSat Feb 17 1996 00:526
    The bottom brush motor in the outside brush station seized up
  because DI water leaked past the bearings.  Everything was wiped
  clean and the motor was replaced.  The motor was recalibrated
  and speed brought into spec.
                                             D shift
47.81IB Upper Arm problemSUBPAC::LANDRYWed Feb 28 1996 19:416
    The system is on WARNING for a broke upper arm in the outside brush
    station.  An order has been placed with OnTrak for this part.  System
    is working fine as is, but the arm does need to be replaced.
    
    						A Shift
    
47.82Spin Station - Wafers Double Stacked & BrokenASDG::POIRIERFri Mar 01 1996 19:316
    
    * Wafers Double Stacked in Spin Station:
      -------------------------------------
      Product wafers broke in the Spin Station Today after being
      double stacked. We reset the Station and no further occurances
      were observed for the remainder of the shift.  
47.83Top roller arm replacedSUBPAC::BJUBINVILLENFD Fire - RescueSat Mar 02 1996 05:3211

           The top section of the front roller arm in the
      outside brush station that was cracked has been replaced.
      
           A little while later, top brush support arm must of
      got bumped during the repair and was causing the wafers to
      ride up on the roller.  It was readjusted and has been ok
      since.

                                    D shift
47.84cracked index station coverSUBPAC::TOOKERTue Mar 12 1996 15:576
    The index station cover is cracked.  I have ordered a replacement
    cover.  It should arrive here by this Friday.  Until, then be very
    gentle with the old one.
    
    Thanks,
    Jim
47.85Send Station ReplacedSUBPAC::BJUBINVILLENFD Fire - RescueSat Mar 23 1996 05:327

           The new SEND station came in and it was replaced.
        Alignments were done and tool passed MQC's.


                                               BJ
47.86Send Station Indexer MotorSUBPAC::LANDRYSat Mar 23 1996 14:1010
The Send Station Indexer would not return to home position after a
cassette of wafers were run.  I could hear the motor trying to drive
elevator home, but there was zero movement.  Traced problem back to
the pulley on the indexer motor.  It was wedged against the motor
retaining bracket.  Adjusted said pulley for proper clearance, reset
up the drive belt tension, and tested several times.  All is working
fine.

      					Doc

47.87roller motor inner stationSTRATA::FINNWed Mar 27 1996 18:3412
	Inner brush station motor roller motor, intermittently failing.
	Removed old motor and shaft and resealed water leak, and changed oring 
        on white plastic seal cap at base of left roller arm. Also found bottom 
        gear missaligned so that it rubbed on bracket, realigned. 
        Reassembled and checked for leaks. Found none. New motor was 
        functional, but required board cal to prevent speed alarms due to high 
        speed.  Ran test wafers aok , returned to manufacturing for  mqc's.


						A shift and HAWK

47.88CMP-SCRUB.A2 shutdown completed for facilities tie-inSUBPAC::BJUBINVILLENFD Fire - RescueThu Mar 28 1996 06:1628

  SHUTDOWN: 5:00 AM ON 3/28/96
  STARTUP:  9:00 PM ON 3/28/96


                CMP-SCRUB.A2 SHUTDOWN CHECKSHEET
                --------------------------------

                                                        Tech.       Date
                                                       -------     -------

  1.  Change RTC status to SCHEDULED DOWN.               LEON      3/28/96

  2.  Remove all PVA scrub brushes and submerge
      in DI water.  If possible, DO NOT remove
      the brushes from their cores.                      LEON      3/28/96

  3.  Perform a complete wipe down and dry of
      all wet surfaces.                                  LEON      3/28/96

  4.  Close main incoming DI Water valve to each
      system and open the DI Water return valve
      (UDR).  This will allow all incoming DI Water
      to flow to UDR.                                    LEON      3/28/96

  5.  Power down each system.                          
                                                         LEON      3/28/96
47.89Inside & outside DI water flow regulatorsFABSIX::R_GEETue Apr 02 1996 07:474
	Found the flow regulators for the inside and outside
	brush stations closed. Set all four flow regulators 
	to the setting of scrub.a1
					B shift
47.90door latch adjustmentFABSIX::L_SHKOLNIKOVWed Apr 03 1996 08:244
           Numerous times system was alarmed for receiving door latch failure
       Door latch has been adjusted. Run 16 wafers - OK!

                                                 D shift  
47.91Inside brush station leakFABSIX::L_SHKOLNIKOVFri Apr 05 1996 07:429
        Inside brush station water leaking through the rear roller shaft
	and through the front arm assembly as well. I took it apart resealed
        water leak (as much as possible without changing roller shaft
        assembly - we don't have it in stock) and changed o-rings.The water
        leak stopped through the front arm but is still coming through the rear arm
        shaft which need to be replaced. I will speak to Jim T. about
        getting this part rushed in.

                                                Leon
47.92Yet another leakSUBPAC::BJUBINVILLENFD Fire - RescueSat Apr 06 1996 06:449
 
           Tools still had a leak coming from the rear
       roller assembly.  The o-ring for the collar didn't
       get replaced yesterday.  Removed rear top roller 
       assembly in the inside brush station and replaced 
       the o-ring inside the white collar.  So far, no leaks...

                                            Leon
47.93Arm replacedSUBPAC::BJUBINVILLENFD Fire - RescueMon Apr 08 1996 16:586
    
    
        Replaced top rear roller arm in the outside 
     brush station.
    
                                       A shift
47.94Burnt out Quartz LampFABSIX::R_POIRIERSun Apr 14 1996 07:577
	The Quartz Lamp in the Spin station burnt out. We replaced
        and gave Scrubber back to Operations.


					B Shift

47.95scheduled contractor work performedSUBPAC::LANDRYTue Apr 23 1996 19:5320
      Today...........

      The DI water input to the room was shut down to perform
      some scheduled work with the DI line AND the OCP's had
      work done concerning the communications/electronics
      package.  This necessitated us to remove ALL the scrub 
      brushes so that they may be placed in water to keep wet.  
      All brushes were labeled and re-installed in their original 
      station/position.

      Ashland performed some electronic/communication work on the
      OCP's.  They verified that the work done to assure proper
      communication between the delivery system/OCP/scrubber.

      Once the water came back on and the system were allowed to
      be started up again, the brushes were installed, test wafers
      were run through the system, and MQC checks were performed.
      The system came back up with no problems.

      						A Shift / Jim T 
47.96Motors offsetSTRATA::RIDLONEliminate the obviousMon May 13 1996 07:487
    The wafer rotation motors (good a name as any) were relocated. Orings
    are being used in the interim until the belts, which have been ordered,
    arrive. The belt is the same that raises and lowers the indexer.
    
    
    					B shift
    
47.97passdownSTRATA::FINNMon May 13 1996 20:1810
	System was not staying in "start mode", Connector J9 had loose pins.
        Also had intermittent problem were belts would not turn on at cassette.
        And intermittent problem with spin station time out error. Adjusted 
        Roller arms at inside brush station.

	Took some time to trace out the OCP/leak EMO relay/power supply 
        circuit. System now up and running.

				A shift/ & "Jim the Toolman Tooker"
47.98New Ammonia Flow Meters installedSUBPAC::LANDRYTue May 21 1996 14:5413
    Whilst doing the Weekly PM today.............
    
    	New flow meters were installed for ammonia use.  The flow meters
    installed dispense a smaller volume than the previously used flow
    meters.  The newer ones were installed to complete the ammonia drip
    manifold project, which was started last week.  Both the Inside and
    Outside brush stations are now complete.
    
    After installation, the flow meters were set dispense an adequate
    amount of 2% ammonia.  
    
    				Respectfully........ the A shift
    
47.99New brush heightsSUBPAC::LANDRYTue May 28 1996 20:2915
    
      Today, we started the system up from a shutdown situation.  At this
      time, new brushes were installed.  The wafer count was zero'd out
      in conjunction with the new brush install.  MQC's passed on the 1st
      pass.

      The New Brush Heights are:  IB-front 570      IB-rear 535
                                  OB-front 620      OB-rear 615

      The new brush heights have been  entered into the latest Weekly PM
      checksheet.  The Weekly PM has been performed for this week.

      						A shift
                                                               
47.100Roller bearing shot!SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringWed May 29 1996 08:327
 
        The rear roller bearing in the send station is cooked.
     None in stock or not a stocked item.  Waiting for the ok
     to "borrow" the set-up from the Wet-Lab.

                           BJ
47.101Roller Bearing Assembly OrderedASDG::POIRIERWed May 29 1996 20:305
    
    Send Station Roller Bearing Assembly:
    ------------------------------------
    
    The assembly has been ordered and will arrive tomorrow at 10:30 a.m.
47.102Send Station Sensor Temporarally ReadjustedASDG::POIRIERThu May 30 1996 20:457
    
    Send Station Sensor Adjusted:
    ----------------------------
    The Send Station sensor has been adjusted to allow the station
    to feed wafers with only one set of belts. We will readjust the 
    sensor when the remaining parts of the pully assembly arrive tomorrow.
    
47.103send st pulley assembly.FABSIX::L_SHKOLNIKOVSat Jun 01 1996 06:157
    
    Send Station Rear pulley assembly 
    --------------------------------
    finally has been installed. The system has performed numerous
    (all night) runs without any further trouble . 
    
                                                d-shift
47.104Megasonics arm home position amplifierFABSIX::R_GEEMon Jun 17 1996 08:157
	Errored for "megasonics arm not home". The arm was in the 
	home position but the sensor wasn't seeing it. The LED was
	out on the photo amplifier. Removed the amplifier and found 
	some contamination on one of the pins. Cleaned and recrimped
	the pins. Re-installed and tested, all set. 

					B shift
47.105You can start_me_up!SUBPAC::LANDRYSun Jun 23 1996 14:1318
Had a problem this morning with the system NOT going into the start
mode.  Rebooted the system, all stations initialized as should, but
when attempting to put the system into start, there was no success.
The system was acting as if an inhibit switch was not being made.  
There is no I/O or other page function to quickly look at to determine
if in fact this was the problem.  Therefore, checked out all the
inhibit switches for proper operation.  Both the IB and OB station
covers and the Spin Station door were checked.  All had 24 VDC
coming to them as should.  Ohm'd out the magnetic switches.  All showed
they were working fine (the meter read OPEN when the switch was not
present and read less than a Ohm when the switch was made).  Went into
the electronics cabinet and checked the pins/connector for the OCP
interface.  The pins appeared to be not making proper contact.  So,
I assured they were making contact by plugging the pins together minus
the connector.  Lo and behold, the system now starts.  Put the con-
nector back together and retried results.  Voila!

      						Dr.
47.106panel / gauges installedSUBPAC::LANDRYTue Jun 25 1996 20:4217
      While the room was shut down for a couple of hours today, we
      took advantage of that time and installed two inline gauges.  
      One for the DI Water and the other for Ammonia.  The gauges
      are mounted above the drip pan to the rear of the machine.
      It is necessary to remove the rear panel (on the Send Station
      end) to see the gauges.  These are currently only for convien-
      ience as they tie into a bigger project, which is alarm warnings
      at the touch screen, which we don't presently have the capability
      of on the systems today.  New software and boards will be installed 
      in the near future which will give us that capability.

      So, the two new gauges have been plumbed in place and are reading
      the pressures.  Be aware that 2% ammonia is flowing to it's new 
      gauge.  The lines were checked for leaks, none were found.

      						A  /  Jim T 
47.107OnTrak after hours supportSUBPAC::TOOKERFri Jun 28 1996 11:306
Here is a telephone number that can be used for contacting OnTrak Systems
field service and technical support after hours.  This number will connect
you to an answering service.  They will forward your message to the people
who are on call.  The number is:  1-888-ONTRAK1

Jim T.
47.108power up o.k.LUDWIG::SWENSONFri Jul 05 1996 20:412
    powered tool up with no issues.
    waiting for di water.
47.109Outside brush station lower brushFABSIX::R_GEETue Jul 09 1996 06:053
	The outside lower brush was hesitating. Had to replace all
	4 core bearings.
					B shift
47.110OB station exit belt rollersFABSIX::R_GEESun Jul 21 1996 08:136
	The wafer wasn't making it to the inside brush station fast
	enough. 
	Had to clean the roller shafts, replace the bearings and a 
	"c" clip.
				
				B shift
47.111ocp mod'sFABSIX::B_FINNSun Jul 28 1996 20:396
	Tied Scrubber A2 into OCP circuit for Scrubber A1. Both these 
        systems are running off one OCP. Tested aok. Ran wires from 
        OCP # 2 to Cmp-Scrub.A3 and labeled for OnTrak rep.

					A shift
47.112Do not use the rubber boots on the carriersSUBPAC::TOOKERFri Aug 16 1996 15:5912
Yesterday during Strasbaugh's visit, we learned that they no longer 
recommend using the rubber boots on the carriers.  Their data now
shows that the boot is interfering with the free movement of the
gyro bearings and therefore impacting uniformity.  The boot was
designed to keep slurry from getting into the gyro bearings.  Be-
cause the carriers are disassembled and cleaned every few hundred
wafers, not using the boot should not cause a problem.  I have
gone through the cmp and carrier build room and removed all of the
boots.

Thanks,
Jim Tooker
47.113Spin station motor controller replacedSUBPAC::TOOKERWed Sep 04 1996 14:4512
CMP-SCRUB.A2 went down last night on D shift because the spin station motor
was not working.  Leon replaced the motor controller PCB.  At the start of
C shift, the motor was working, but needed to be calibrated.  The motor
controller PCB comes with a manual from the OEM supplier that has a calibration
procedure.  Please do not attempt to use this procedure.  The correct
procedure can be found in the OnTrak service manual.  We performed the
calibration as per the manual and the system is now operating correctly.

Also, as a separate item, we made a slight adjustment to the spin lock
solenoid.

                                         Jim T./Jim S./John P.
47.114Adjusted spin station wafer sensorSUBPAC::TOOKERWed Sep 04 1996 16:414
    Adjusted spin station wafer sensor amplifier and processed 25 wafers
    to verify.
    
    Jim T.
47.115SPin Station Motor COntrollerASDG::POIRIERWed Sep 04 1996 20:469
    
    Spin Station Motor Controller:
    -----------------------------
    We had trouble getting the span pot to yield 1800 rpm. A jumper was
    moved on the board and the span pot responded correctly. Both the zero
    and span pots were adjusted to bring the spin station rotation into
    spec. 
                                          C Shift.
    
47.116robot fingers not closed.FABSIX::L_SHKOLNIKOVThu Sep 05 1996 08:3114
    	The system robot functions were messed up this morning.  The
    fingers were not closing to their proper position, causing the system
    to "timeout" Robot fingers not closed. Found the several problems with the
    system.  Problem one was the flags for the open + closed potitions were 
    not being made.
    After this was little adjastment and, we attempted to run some wafers. 
    We found the gripper arms to be overtraveling,causing the grippers to flex
    back. The closed position was readjasted. Transfered 75 wafers - robot
    works great but occasionally, error lost wafer at spin comes up. 
    So here we are now.

    
                                                         D shift
47.117High ParticlesASDG::POIRIERThu Sep 05 1996 20:259
    
    Particle Problems:
    -----------------
    The system failed MQC's at 6:00 p.m. for high particles. (500+)
    Three cassettes of wafers have been run thru the tool in an attempt
    to purge it clean. D Shift, run MQC's when you get in.
    
                                                          C Shift.
    
47.118Receive station door lockFABSIX::R_GEETue Sep 10 1996 08:386
	The system was alarming for receive station door lock
	failure. Found the door lock wasn't retracting back into
	the air cylinder fast enough. Cleaned out the air cylinder
	venting port and adjusted the door lock assembly, all set.

				B shift
47.119O.B. Tach problemFABSIX::R_POIRIERSun Sep 22 1996 07:406
    System was alarming for Outside brush station out of spec.
    We found a a loose wire going to the Tach. Re-soldered the
    wire back and problem went away.
    
    						B Shift
    
47.120Robot finger issueSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Sep 27 1996 22:1310

      Came in with tool down because the robot fingers were
   failing to close properly...

      Powered down tool with no success.  We ended up bringing
   in the finger motor count by 10.  Tool ran 25 wafer verify 
   with no problems.
                                       BJ
           
47.121OnTrak upgrade work continuesSUBPAC::TOOKERThu Oct 03 1996 17:066
    OnTrak field service continued work on the pressure interlock upgrades.
    New software was installed and all electronics were installed.  He
    is lacking some plumbing fittings that should be here tomorrow.  When
    OnTrak leaves for the day, the tool will be available for Operations.
    
    Jim
47.122UPGRADESSTRATA::WESTPHALThu Oct 03 1996 20:285
SCRUB.2 - Ontrack Rep was in to do the same upgrades that were done yesterday 
	  on SCRUB.1.  Rep installed sensors to monitor various pressures and 
	  check for a heat lamp burn out condition.  He will return tomorrow to
	  install the software on both tools.
 
47.123Update on pressure interlock upgradesSUBPAC::TOOKERFri Oct 04 1996 17:0415
Today OnTrak field service continued work on the pressure interlock upgrades.
They have now completed all necessary work on cmp-scrub.a1/.a2/.a3 except
for the following open items:

  1) CMP-SCRUB.A1:  Needs new software installed to enable the new pressure
          &         interlocks.  The new software will arrive on Monday and
     CMP-SCRUB.A2   OnTrak will be back on Tuesday to install.

  2) CMP-SCRUB.A3:  Needs final connection of the NH4OH pressure interlock
                    to the incoming NH4OH supply line.  We will do this
                    ourselves because all three scrubbers will need to go
                    down for 45 minutes to perform this task.


    Jim
47.124software / alarms tied inSUBPAC::LANDRYTue Oct 08 1996 20:369
      New software was installed today.  The new software has allowed
      for the DI Water and Ammonia to be tied into the alarm package.
      Both fluids were tested to assure that the alarms worked when
      a lack of fluid pressure would occur.  The software also ties
      in the Heat lamp.  If the lamp fails to illuminate, the alarm
      sounds.  This featiures was tested as well.

      					A shift
47.125IGNORE THIS ALARMLUDWIG::WESTPHALThu Oct 10 1996 20:355
SCRUB.A2 - Both scrubbers have been alarming been alarming intermittently for a 
	  "leak detected in machine floor".  The rep believes the problem is 
	   being caused by a floating ground.  This does not effect the proper
	   operation of the tool.  If the error is received just clear and 
	   continue running.  ONTRAK is working on a fix. 
47.126OCP alarmFABSIX::B_FINNTue Oct 15 1996 15:426
	OCP alarm for NH4 no flow. Flow was good, tried to reset OCP,
        but could not, had to emo and reboot OCP controller


					A shift
47.127pulley slippingFABSIX::B_FINNSun Nov 10 1996 19:2511
	Scrub.A2's outer station transport belts failed to come on. 
Inspection showed the motor and primary belts turning but secondary 
pulley slipping. All set screws on the pulley/shaft were tight. 
Removed pulley and found its inner bushing slipping within the pulley. 
Non in stock. We used a center punch to drive the bushing out. 
Applied "locktight" and reinstalled bushing to pulley, then reinstalled 
pulley on shaft. Tested aok


					A SHIFT  and Jim Sadin
47.128Low Ammonia alarm problem solvedSUBPAC::LANDRYMon Nov 18 1996 19:197
After a couple of days of intermittent "low ammonia" alarms, Ashland
was able to find the fault.  They traced the problem to a stuck valve
within their control panel down stairs.  The problem was fixed before
10:00 AM, with no alarms happening the rest of the shift.  Hopefully,
it stays that way.

                                  A shift
47.129Glass Bearings Replaced On the Outside Brush Station.ASDG::POIRIERThu Nov 21 1996 19:3210
    
    Outside (Top) Brush Rotation Problems:
    -------------------------------------
    The system errored for "brush rotation rpm" on the outside brush
    station today. All four glass bearings were replaced on the top brush 
    and the problem ceased. 
    
    
                                                          C Shift.
    
47.130Brush motor replacedSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Nov 23 1996 05:189
          Tool kept alarming for outside top motor speed
      out of spec.  Tweaked the speed to match the output
      and matched with tach.   It ran fine for several hours
      then it dragged low again.  This time, the motor was
      replaced and checked out ok...

                                      D shift
           
47.131cassettes brokenFABSIX::M_RIDLONEliminate the obliviousWed Dec 04 1996 03:598
    Timed out for robot not leaving spin station. Found the cassette switch
    not being made because the cassette had a broken tab. Threw it away and
    replaced with a whole cassette. Happened again later in the night and
    again it was the cassette which again it had a broken tab. Instructed
    the operator on how to check for this.
    
    		
    					B-shift
47.132IB & OB HANDLING PROBLEMSFABSIX::B_WESTPHALWed Dec 11 1996 19:248
SCRUB.2 - Down for upper brush on the IB station not going down and other
	  general handling problems.  The J.P & J.S. team corrected the Brush
	  not going down problem by adjusting the brush's lift cylinder.  The 
	  rollers for both the IB & OB stations were changed out and adjusted. 
	  A Transfer belt in  the IB station was also found to be defective and
	  was replaced.

47.133TOP ROLLER ARM REPLACEDSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Dec 12 1996 06:137
    
    
          Replaced the front top roller arm in the inside brush station.
       It was observed to having a hair-line crack during the Quarterly
       PM. 
                                          D shift
    
47.134Recieve Station Lifter Failed To Rise to the top PositionASDG::POIRIERSat Jan 18 1997 19:3427
    
    Recieve Station Lifter:                                       
    ______________________
    
    The Lifter in the Recieve Station failed to rise to the top position
    today. This caused the Robot to drop a wafer into the base of the tool.
    The tool was powered down and rebooted. The lifter then worked
    normally. This condition has occured a few times over the last month
    on C shift. If it happens again, we might want to replace the pia board
    for the Recieve Station. Ontrak talked about this type of intermittent
    symtom in class. The usual cause is a chip on the pia board. We kept
    the tool up due to the high usage in the area today. 
    
    FYI: Except for the time lost due to the breakage on A2 this morning, all
         five polishers cranked for the entire shift. The only issue
         affecting thruput was the Operators waiting at the UV1050's 
         and at the Ontrak Scrubbers.  
    
         If we put B4 on line and have a day like we did today, the
         effect of adding this tool will be less than optimal.
     
                                        (Another Scrubber & UV1050?)
                                        (Are they comming soon?) 
    
    
                                                              C Shift.
                                                          
47.135bad load station roller bearingFABSIX::J_SADINFreedom isn&#039;t free.Sun Jan 19 1997 18:2711
    
    
    	Load Station Transport roller bearing and o-ring:
    	------------------------------------------------
    
    
    	Rear load station rollers were found not to be spinning. Replaced
    left roller bearing (it was rusted/frozen) and large o-ring to rear
    load-station rollers (old o-ring was badly stretched). 
    
    
47.136dropped wafer in unload station againFABSIX::J_SADINFreedom isn&#039;t free.Tue Jan 21 1997 18:0218
    
    
    Receive station lifter (again):
    
    
    	The lifter (blade) in the receive station didn't come to its top
    position and dropped a wafer that was in the process of unloading.
    Rescued wafer from the bottom of the tool (didn't break). Reset unload
    station ok (blade came back to top position), but then machine robot
    wouldn't unload wafer from spin station. Had to power down the tool and
    restart to get the wafers unloaded. All seemed ok after that. 
    
    	As John mentioned in .134, this has happened before, possibly due
    to a faulty chip in the pia board. We'll check and see if there is one
    in stock. Tool is back to production for now (lotsa work in the area).
    
    	A-shift
    
47.137status of software upgradeSUBPAC::TOOKERWed Jan 22 1997 15:418
    OnTrak field service came in today to install a new revision of
    software.  This new revision will add a small time delay before
    alarming for low DI Water or NH4OH pressure.  This will help to
    eliminate alarms due to pressure fluctuations from valves actuating.
    However after doing the installation, OnTrak found a bug in the
    software.  He removed the new rev. and re-installed the old rev.
    
    Jim T.
47.138Software Upgrade - UnsucessfulASDG::POIRIERWed Jan 22 1997 19:3311
    
    Software Upgrade Unsucessful:
    ----------------------------
    The OnTrak rep arrived with the new version of software. This version 
    was installed to prevent the eronious Ammonium Hydroxide and D.I.
    Water "Low Pressure" Alarms. The system would not detect any pressure
    after this latest version was installed. The Ontrak software Dept is
    going to make another attempt at getting this problem solved. 
    
                                                             C Shift.
    
47.139robot to blade transfer problemSUBPAC::LANDRYSun Feb 23 1997 19:3832
    Whilst the Sunday morning EE passdown meeting was taking place, the
    system's robot to blade transfer messed up, causing the wafer to drop
    to the base, and break upon impact.  Previous occurances of this have
    happened in the past week or two (with A shift seeing it only once
    prior).  Seeing no one actually witnessed this mis-hand-off (or any of
    the other occurrnces to our knowledge), it was being "assumed" that the
    blade didn't come up to it's proper height.  We on A shift can not
    confirm if any other shift actually knows if the blade was low or not.
    So, we sat and watched the system run for awhile.  The cover over the
    blade mechanism was removed.  The motor, gears, belt, and sensors were
    all running properly.  No problems were exhibited.  So, we decided to
    put the cover back on and close up the electronics cabinet.  Once the
    electronics door was closed, we rebooted the system and resumed
    testing.  Please note that we left the few wafers in the unload station
    in the cassette before resuming.  As it was lifting the wafers up to
    detect an occupied slot, it missed picking one up, causing the blade to
    rise between two wafers.  We immediately halted and removed the
    cassette from the unload station.  At this moment, we noticed a great
    deal of slop in the cassette.  The cassette was checked and we found
    that a locating tab was sheared off of the cassette.  This caused the
    cassette to rotate in the unload station (which vibration of the unload
    handler movement caused).  This caused the blade to cassette slotting
    to be off, which in turn caused the wafer to make contact with a slot
    guide witin the cassette, hence sending the wafer down to destruction.
    After we installed a cassette of the utmost craftsmanship, we could not
    duplicate the robot to blade mishap.
    
    After this monumental discovery, we purged several cassettes from the
    fab which had this broken / or missing locating tab.  
    
                                                      A shift
    
47.140lifter not highFABSIX::M_RIDLONEliminate the obliviousMon Feb 24 1997 03:1119
    Tool dropped a product wafer at around 10pm. The wafer didn't break and
    was inspected for chips (none). The lifter was observed as not coming up
    to the proper height. We checked the cassettes as well and found them
    to be in satifactory condition. Pressing the RESET STATION didn't correct
    the lifter, however, pressing RESET RECV did. This problem seems to be 
    extremely intermittent. We decided to swap the intelligent stepper
    board that drives the lifter motor (also drives the receive motor) with
    the board that drives the send motor (other channel is spare). We then
    ran 150 wafers with no alarms or dropped wafers. (The spare channel and
    the lifter channel were the same between the boards for those inquiring 
    minds). The intelligent stepper boards are the 4 boards next to the DC
    motor control boards. We swapped the 2nd and 3rd boards (ribbon cables 
    are clearly marked). We left them swapped as this is too intermittent
    to determine if this is the cause. The part number for the board is 
    28-0075-001.
    
                				B-shift
    
    				 
47.141Knife belt/bearings replacedSTRATA::FINNSun Mar 02 1997 19:1811
	Dropped a wafer during unload at receive station knife. Removed old 
Knife Belt and inspected. Replaced the belt at EE's directive. Also replaced 
two bearings in the belts upper sprocket assembly. Adjusted the new belt's 
tension several times to insure it was not skipping teeth and that it's 
travel was repeatable.

Ran 25 wafers aok and returned system to manufacturing.


					A shift
47.142broken wafer in spin stationFABSIX::J_SADINFreedom isn&#039;t free.Wed Mar 12 1997 19:337
    
    
    	wafer pulverised in spin station at full rpm. cleaned tool and
    running 25 dummy wafers. Leaving for D-shift to verify fix and put tool
    up. 
    
    	a-shift
47.143replaced o-ring on outside brush stationFABSIX::J_SADINFreedom isn&#039;t free.Mon Mar 17 1997 19:358
    
    
    	replaced o-ring 202-113 that drives the incoming transport rollers
    on the outside brush station (rollers were not spinning). running well
    now.
    
    
    	a-shift
47.144o-ring replacedFABSIX::J_SADINFreedom isn&#039;t free.Sun Mar 23 1997 12:259
    
    
    	replaced incoming transport roller o-ring again (on outside
    brush station). Rollers were not spinning properly. Appears to be
    working properly now.
    
    
    a-shift
    
47.145Wafer left at SPINSUBPAC::BJUBINVILLECMP Equipment EngineeringSat Mar 29 1997 04:166
    
          A wafer was left in the spin nest after the robot
       either missed it or never picked it up.  Ran a 25 wafer
       VERIFY without any other problems.
    
                                                   D shift
47.146wafer roller drive beltFABSIX::J_SADINFreedom isn&#039;t free.Sun Mar 30 1997 19:5612
    
    
    
    	The drive belt for the front side wafer roller (OB station) fell
    off today. Adjusted the pulleys for the belt so that the belt would not
    ride up so close to the top of the back pulley (allowing it to slip
    off). Has been running fine since.
    
    
    	a-shift
    
    
47.148Sensor not seeing these particular wafersFABSIX::B_WESTPHALFri Apr 04 1997 20:0312
    
CMP-SCRUB.A2 - Down for waiting for wafers in the spin station.  Looked on the 
	I/O screen and seen that the wafers were not being sensed in the spin
	station.  The wafers were removed from the tool and sent through
	SCRUB.A3  while looking at problem on SCRUB.A2.  SCRUB.A3 then alarmed 
	for the same thing.  Found out that the lot (PE0028) was an engineering
	lot.  Seen no reason to adjust the sensor to these particular wafers 
	so I pushed the last 10 Engineering wafers through on A3 & ran a boat 
	on A2 with no errors.  Gave the tool back to production.



47.149Robot Stopped At Spin Station / Tool Locked Up.ASDG::POIRIERThu Apr 10 1997 20:3210
    
    Robot Errored AT Spin Station:
    -----------------------------
    The robot stopped while attempting to pick up a wafer at the Spin
    Station. The system locked up after the robot stopped. The tool
    was powered down and rebooted. There have been no reoccurences 
    of this problem since 10:00 a.m. this morning.
    
                                                           C Shift.
     
47.150Robot dropping wafersSUBPAC::BJUBINVILLECMP Equipment EngineeringSat Apr 12 1997 07:0212
          
          The robot missed/dropped a wafer the spin station.  25
      wafers were run and there weren't any problems.  Later on,
      it did it again.  The handling was checked and the only thing 
      that was found was that the robot was gripping the wafers too 
      tight.  You are suppose to be able to rotate the wafers when the 
      robot fingers are closed and I couldn't even move it a 1/2".
      I backed off on the off-sets and ran more wafers.  The problem
      hasn't come back but I'm a little unsure if that was the real
      fix.

               
47.151Wafers dropped by robotFABSIX::R_POIRIERMon Apr 14 1997 08:4123
		Scrubber was dropping wafers in spin station.
		This happened three times. The first two times
		we were able to recover because fingers knew that
		a wafer wasn't present, but the last time the
		robot thought it had the wafer. Robot tried
		to put wafer into lifter but no wafer was
		present. It had fallen into the spin station.
		While trying to recover, open fingers flag 
		didn't trip Open fingers sensor and the cam 
		continued to turn until flags were 180 degrees
		out. The fingers would open when commanded to 
		close and close when commanded to open. We 
		re-adjusted the flags to take care of that problem
		and then tightened the tension of the fingers
		using Teach. (Increased from 130-150)
		

					B Shift



		
47.152Heat lamp blownFABSIX::R_POIRIERSun Apr 27 1997 07:549

		Scrub A2 blew it's heat lamp last nite.
		Found a spare in the vidmar and replaced
		bad lamp.


				B Shift

47.153Leaking shaftsFABSIX::O_WHITFORDLead, follow or step aside.Sun Apr 27 1997 20:0212
    During the weekly, it was noted that there was some pretty considerable
    leakage from both the outer and inner brush stations. Working under
    some time constraints, only the outer station leak was addressed at this
    time. Found the front roller shaft was leaking. When it was removed, we
    found the upper seal was installed up-side down (spring facing up), 
    the outer face was missing a chunk and scored. We replaced the shaft
    and managed to crack the lower support are while remounting it to the
    shaft. Replaced that, lined it up and ran some verification wafers with
    no problems, check particles and they were in spec. Will try to
    schedule time tomorrow AM to work the inner brush station leak issue.
    
                                        Dave W. 
47.154Completion of shaft replacementsFABSIX::O_WHITFORDLead, follow or step aside.Mon Apr 28 1997 20:2410
    Replaced the leaking shaft on the inner brush station today. Follow up
    visit revealed that the leak(s) have been slowed, but there is still a
    persistant leaker somewhere at the rear of the tool in the vicinity of
    the inner brush station. Re checked all water line connection, all
    tight and no evidence of leakage from one of these. No evidence of more
    leaking shafts. ?? THAT much accumilation from between the out and
    ineer brush station? Possible but doesn't seem there would be THAT much 
    in only a couple hours of operation..... 
    
                                                Dave W
47.155TOOL LOCKED UPLUDWIG::STANDINGThu May 01 1997 20:164
    SCRUB.A2   Went down for lockup, with wafers stuck in the robot arm and
    both scrub stations.   EMOd tool and recovered wafers.  I think that
    the unload cassette came off the cassette switch.  Returned tool to
    manufacturing.
47.156Lifter bearingsFABSIX::O_WHITFORDLead, follow or step aside.Sun May 11 1997 11:3910
    Picked up from B shift.
    
        Looked over lifter mechanism and found top sprocket shaft tilted
    at a crazy angle. Lowered lifter all the way down and removed support
    housing for the top sprocket shaft. Left end bearing was falling apart.
    
        Replaced bearings and ran verification wafers with no problems.
    
                                            A shift
    
47.157Current brush heights as of 5/13/97SUBPAC::BJUBINVILLECMP Equipment EngineeringTue May 13 1997 22:478
    
             The below brush height readings are the current
          ones on the tool and have been verified.  Please make
          note of them in the next weekly PM.
    
                            Front     Rear
                 Inside     .395      .385
                Outside     .375      .332
47.158WW46 Weekly PM completedFABSIX::R_POIRIERSat May 17 1997 22:4539
--------------------------------------------------------------------------------
        OnTrak DSS-200 Scrubber Maintenance Procedures HF6O-0170
           Weekly Preventative Maintenance Check Sheet, Rev D

 DATE: 05/12/97  SYSTEM RTC CODE: CMP-SCRUB.A2  TECH. NAME:bd=Bruce Durgin
                                                           

  6.1.1  Check the CDA pressure for 50 PSI		Pre-Adjust: 50

  6.1.2  Check the D.I. Water pressure for 40 PSI       Pre-Adjust: 40

  6.1.3  Check the Drip Pan for contamination           Completed:  bd

  6.1.4  Check all tubing/fittings for leaks		Completed:  bd

  6.1.5  Check all wafer transport belts		Completed:  bd

  6.1.6  Clean the Send module and Scrub modules  	Completed:  bd

  6.1.7  Check scrub brushes for contamination		Completed:  bd

  6.1.8  Check scrub brush height in both modules	Completed:  bd

                   Brush heights

                         OUTER

                            Front    .375
                            Rear     .335

                        INNER 

                            Front    .405
                            Rear     .422

       


47.159Wafer lost in the receive?FABSIX::B_WESTPHALBradley BP#661 EXT# 6610Thu May 29 1997 20:434
CMP-SCRUB.A2 - Down for losing a wafer in the receive station. No definite fix
	was found.  2 boats of wafers were run without repeating the problem 
	returned the tool to operations.